Circular No. 6256

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995al IN NGC 3021 J.-y. Wei, W.-d. Li, Y.-l. Qiu, and J.-y. Hu, Beijing Astronomical

Observatory, report that a spectrogram (range 300–900 nm, resolution 19.5 nm/mm), taken on Nov. 3.88 UT with the 2.16-m telescope at Xinglong station, shows that this object is a type-Ia supernova near maximum. The spectrum shows conspicuous Si II 635.5-nm absorption at 615 nm and resembles the spectrum of the type-Ia SN 1981B near its maximum on 1981 Mar. 6–9. Other distinct absorption features in the spectrum include S II (545.1 nm), Si II (506 nm), and Mg II (426 nm).

K. Ayani, Bisei Astronomical Observatory (BAO), Okayama; and H. Kawakita, Nara, Japan, report that a spectrogram (range 380–730 nm, resolution 1.3 nm; 20-min exposure) of SN 1995al obtained on Nov. 4.77 UT with the BAO 1.0-m telescope confirms the results obtained by the Beijing group: “Absorption features of Si II (635.5 nm), S II (545.4 nm), and Mg II (448.1 nm) and complex absorptions of Si II and S II short of 500 nm are clearly seen. Balmer lines are absent, and the Si II 597.2-nm absorption line is still weak, indicating that this is type-Ia supernova at maximum; this spectrum resembles that of SN 1981B at maximum (Branch et al. 1983, Ap.J. 270, 173). The Si II 635.5-nm absorption line is blueshifted with respect to the galaxy (NGC 3021, recession velocity 1500 km/s) by 14 000 km/s and has a FWHM of 6900 km/s.”

T. Barker, Wheaton College, reports that a video recording of two unfiltered 35-s CCD images on Oct. 24.40 UT shows SN 1995al to be ∼ 1 mag fainter than the foreground star (V ∼ 16.0) reported by Pollas on IAUC 6255. On Oct. 27.40, the same equipment and exposure times show the supernova ∼ 0.1 mag brighter than this same star.

Magnitude estimates (visual unless otherwise noted): Nov. 3.04 UT, 13.0 (S. Moretti and G. Contoli, Lugo di Romagna, Italy; unfiltered CCD); 3.19, 12.8 (B. H. Granslo, Fjellhamar, Norway); 4.14, 13.1 (T. Vanmunster, Landen, Belgium); 4.20, 12.7 (Granslo); 5.11, 13.0 (P. Schmeer, Bischmisheim, Germany); 6.19, 13.0 (Schmeer).

NOVA CASSIOPEIAE 1995 Visual magnitude estimates: Sept. 20.87 UT, 8.9 (E. Broens, Mol,

Belgium); Oct. 1.18, 9.1 (Granslo); 16.96, 8.9 (A. Pereira, Cabo da Roca, Portugal); 24.12, 9.2 (C. E. Spratt, Victoria, BC); Nov. 3.07, 8.5 (Granslo); 4.77, 8.4 (M. V. Zanotta, Laino, Italy); 5.861, 8.4 (Schmeer).

1995 November 6 Daniel W. E. Green